OpenStudy (anonymous):

The time t required to drive a certain distance varies inversely with the speed r. If it takes 12 hours to drive the distance at 60 miles per hour, how long will it take to drive the same distance at 85 miles per hour?

OpenStudy (anonymous):

@SolomonZelman @ganeshie8 @abb0t

OpenStudy (anonymous):

@Hero @myininaya

OpenStudy (anonymous):

\[\frac{ 60 }{ 12 }=\frac{ 85 }{ t }\] i believe you have to set up a proportion like this then solve

OpenStudy (anonymous):

@elementwielder well t=17?

OpenStudy (anonymous):

yes that is the answer

OpenStudy (kendricklamar2014):

This is how you solve this question: \[12 * 60 = 720\] \[\frac{ 720 }{ 85 } = 8.47...\]